A metal bar at 50cm long at 15°C is heated to 85°C. If it expands by 0.088cm ,determine it's linear expansitivity


1
Expert's answer
2021-02-22T10:20:18-0500

By definition, the increase in length will be:


"\Delta L = \beta L_0\Delta T"

where "\beta" is the linear expansivity, "L_0 = 50cm" is the initial length, and "\Delta T = 85\degree C - 15\degree C = 70\degree C" is the change in temperature. Since "\Delta L = 0.088cm", the linear expansivity is:


"\beta = \dfrac{\Delta L}{L_0\Delta T}\\
\beta = \dfrac{0.088}{50\cdot 70} \approx 2.51\times 10^{-5}\degree C^{-1}"

Answer. "2.51\times 10^{-5}\degree C^{-1}".